# Parameter Efficient Continual Learning with Dynamic Low-Rank Adaptation

### v2

@@ B.14 Inference time comparison in Class-IL
L2P [72] 9.53
DualPrompt [70] 9.51
CODA-P [60] 9.47
Adam-ssf [77] 10.03
SEMA [65] 7.39
@@ C.1 Backbones
To demonstrate the versatility of PEARL across different architectures, we also explore three distinct backbones: (i) Convolutional networks such as ResNet-18 and a ResNet-18-like architecture incorporating Blueprint Separable Convolutions (BSC) [43], and (ii) the Vision Transformer (ViT) Base 16 pre-trained on ImageNet-21k.
With regard to the PEARL (R18), the backbone is same as the one used in Mammoth CL benchmark [10]. In line with previous works, the backbone is randomly initialized and trained for CL tasks. However, we introduce PyTorch ModuleLists all throughout to make it easier to add / remove convolutional and BatchNorm layers at will. The ResNet-18 backbone consists of 4 layers with 2 blocks each. Each block in turn consists of two 3x3 convolutional layers with a BatchNorm and ReLU activation in between and after. ResNet-18 also entails shortcut connections in the form of a 3x3 convolutional layer followed by a BatchNorm layer and ReLU activation. Within PEARL (R18), we reserve separate classification and BatchNorm layers for each task. As the backbone is not pre-trained, we treat the first task as the reference task and conduct dynamic allocation from the second task onward. LoRA components are allocated for every convolutional layer. Their rank is dynamically decided after SVD decomposition of respective task vectors. Even after dynamic allocation and further training, LoRA components, BatchNorm layers and a corresponding classification layer are kept separate for each task.
Li et al. [43] introduced blueprint separable convolutions (BSC) to optimize and reduce redundancy in convolutional operations. Essentially, BSC entails a 1x1 operation (point-wise filter) followed by a depth-wise convolution layer. This separability reduces the memory and computational footprint resulting far more efficient model. Such a model is well suited for parameter isolation CL approach as it reduces overall model size when dealing with longer task sequences. PEARL (BSC) backbone consists of 3 layers with 2 blocks each. Within each block, the organization of layers is same as that of ResNet-18 except that traditional convolutional layers are replaced by BSC layers. As the BSC backbone is a custom, small network, pre-trained model is not readily available. Therefore, we treat the first task as the reference task and conduct dynamic allocation from the second task onward. LoRA components are allocated for every point-wise filters as they have the majority of the footprint in terms of number of parameters. Since depth-wise filters have relatively lower footprint, we allocate LoRA components and select the rank of the task vector after decomposition. Same as in ResNet-18, even after dynamic allocation and further training, LoRA components, BatchNorm layers and a corresponding classification layer are kept separate for each task.
@@ C.2 Datasets and settings
We evaluate PEARL across two distinct scenarios of CL: Class-IL and Task-IL. In both these settings, the model encounters a new set of classes in each task and must learn to distinguish all classes encountered thus far after each task. In practice, we split CIFAR-100 [38], TinyImageNet [40], and ImageNet-R [27] into different tasks depending on the number of tasks as in Table 17.
@@ C.3 Training details
For experiments concerning PEARL (ViT) on the ImageNet-R dataset, we use the Adam optimizer [36] with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The batch size is set to 128. The ViT model is fine-tuned on the current task with a learning rate of $5\times 10^{-3}$ for 25 epochs ($E_{1}$). Following this, the LoRA weights are trained separately for an additional 25 epochs ($E_{2}$) using a learning rate of $5\times 10^{-4}$. We observed that a cosine-decaying learning rate schedule outperforms a constant learning rate. Regarding the LoRA parameterization, the rank is determined dynamically based on the criteria outlined in our approach, and the LoRA alpha parameter is set to twice the rank.

### v3

@@ B.14 Inference time comparison in Class-IL
L2P [wang2022learning] 9.53
DualPrompt [wang2022dualprompt] 9.51
CODA-P [smith2023coda] 9.47
Adam-ssf [10.1007/s11263-024-02218-0] 10.03
SEMA [SEMA] 7.39
@@ C.1 Backbones
To demonstrate the versatility of PEARL across different architectures, we also explore three distinct backbones: (i) Convolutional networks such as ResNet-18 and a ResNet-18-like architecture incorporating Blueprint Separable Convolutions (BSC) [li2022blueprint], and (ii) the Vision Transformer (ViT) Base 16 pre-trained on ImageNet-21k.
With regard to the PEARL (R18), the backbone is same as the one used in Mammoth CL benchmark [buzzega2020dark]. In line with previous works, the backbone is randomly initialized and trained for CL tasks. However, we introduce PyTorch ModuleLists all throughout to make it easier to add / remove convolutional and BatchNorm layers at will. The ResNet-18 backbone consists of 4 layers with 2 blocks each. Each block in turn consists of two 3x3 convolutional layers with a BatchNorm and ReLU activation in between and after. ResNet-18 also entails shortcut connections in the form of a 3x3 convolutional layer followed by a BatchNorm layer and ReLU activation. Within PEARL (R18), we reserve separate classification and BatchNorm layers for each task. As the backbone is not pre-trained, we treat the first task as the reference task and conduct dynamic allocation from the second task onward. LoRA components are allocated for every convolutional layer. Their rank is dynamically decided after SVD decomposition of respective task vectors. Even after dynamic allocation and further training, LoRA components, BatchNorm layers and a corresponding classification layer are kept separate for each task.
li2022blueprint introduced blueprint separable convolutions (BSC) to optimize and reduce redundancy in convolutional operations. Essentially, BSC entails a 1x1 operation (point-wise filter) followed by a depth-wise convolution layer. This separability reduces the memory and computational footprint resulting far more efficient model. Such a model is well suited for parameter isolation CL approach as it reduces overall model size when dealing with longer task sequences. PEARL (BSC) backbone consists of 3 layers with 2 blocks each. Within each block, the organization of layers is same as that of ResNet-18 except that traditional convolutional layers are replaced by BSC layers. As the BSC backbone is a custom, small network, pre-trained model is not readily available. Therefore, we treat the first task as the reference task and conduct dynamic allocation from the second task onward. LoRA components are allocated for every point-wise filters as they have the majority of the footprint in terms of number of parameters. Since depth-wise filters have relatively lower footprint, we allocate LoRA components and select the rank of the task vector after decomposition. Same as in ResNet-18, even after dynamic allocation and further training, LoRA components, BatchNorm layers and a corresponding classification layer are kept separate for each task.
@@ C.2 Datasets and settings
We evaluate PEARL across two distinct scenarios of CL: Class-IL and Task-IL. In both these settings, the model encounters a new set of classes in each task and must learn to distinguish all classes encountered thus far after each task. In practice, we split CIFAR-100 [krizhevsky2009learning], TinyImageNet [le2015tiny], and ImageNet-R [hendrycks2021many] into different tasks depending on the number of tasks as in Table 17.
@@ C.3 Training details
For experiments concerning PEARL (ViT) on the ImageNet-R dataset, we use the Adam optimizer [kingma2015adam] with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The batch size is set to 128. The ViT model is fine-tuned on the current task with a learning rate of $5\times 10^{-3}$ for 25 epochs ($E_{1}$). Following this, the LoRA weights are trained separately for an additional 25 epochs ($E_{2}$) using a learning rate of $5\times 10^{-4}$. We observed that a cosine-decaying learning rate schedule outperforms a constant learning rate. Regarding the LoRA parameterization, the rank is determined dynamically based on the criteria outlined in our approach, and the LoRA alpha parameter is set to twice the rank.